OpenStudy (anonymous):

Five added to 2 times a number is equal to 7 less than 4 times the number. What is the number?

OpenStudy (anonymous):

@justsmile531

OpenStudy (anonymous):

@jbas

OpenStudy (anonymous):

\[5+2\times n=7-4n\]

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

I agree with Night

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

they wont let me put in the letters

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

Just type the letter n

OpenStudy (anonymous):

ok wat else

OpenStudy (anonymous):

Just type this: 5+2×n=7−4n

OpenStudy (anonymous):

so its 4n

OpenStudy (anonymous):

No, its this 5+2×n=7−4n. That's the full answers.

OpenStudy (anonymous):

i cant put in all that

OpenStudy (anonymous):

i only can put like 4 numbers in

OpenStudy (anonymous):

Can you take a screenshot of the question?

OpenStudy (anonymous):

Unit 2 | Lesson 8 Lesson Assessment: Variables on Both Sides of an Equation Quiz Question Navigator Enter the answer. 5. Five added to 2 times a number is equal to 7 less than 4 times the number. What is the number?

OpenStudy (anonymous):

A screenshot.

OpenStudy (anonymous):

no i cant do it

OpenStudy (anonymous):

there are no answers its just a box

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

I really need some help guys

OpenStudy (anonymous):

tag me i can try to help you